Find the isogonal trajectories of the family of curves x2+y2=c if θ=45°.

1
Expert's answer
2021-11-26T05:22:02-0500

The given family of curves are

"x^{2}+y^{2}=c"

Differentiating w.r.t. "x" , we get

"2 x+2 y(d y / d x)=0\\

x+y(d y / d x)=0 ...(i)"

Replacing

"\frac{d y}{d x}\ by \ \frac{\frac{d y}{d x}-\tan 45^{\circ}}{1+\frac{d y}{d x} \tan 45^{\circ}}=

\frac{\frac{d y}{d x}-1}{1+\frac{d y}{d x}}"

in Eq. (i), we get

"x+y\left(\frac{\frac{d y}{d x}-1}{1+\frac{d y}{d x}}\right)=0"

 

"\begin{aligned}

&x\left(1+\frac{d y}{d x}\right)+y\left(\frac{d y}{d x}-1\right)=0 \\

&\Rightarrow \quad(x+y) \frac{d y}{d x}=(y-x) \\

&\Rightarrow \quad \frac{d y}{d x}=\frac{y-x}{y+x} \quad \ldots(\mathrm{ii})

\end{aligned}"

which is a homogeneous differential equation.

"\begin{aligned}

&\text { Put } y=v x \Rightarrow \frac{d y}{d x}=v+x \frac{d v}{d x} \\

&\Rightarrow \quad v+x \frac{d v}{d x}=\frac{v-1}{v+1} \\

&\Rightarrow \quad x \frac{d v}{d x}=\frac{v-1}{v+1}-v=\frac{v-1-v^{2}-v}{v+1} \\

&\Rightarrow \quad\left(\frac{v+1}{v^{2}+1}\right) d v=-\frac{d x}{x}

\end{aligned}"

Integrating, we get

"\begin{aligned}

&\frac{1}{2} \log \left|v^{2}+1\right|+\tan ^{-1} v=c_1-\log x \\

&\Rightarrow \quad \log \left|x^{2}+y^{2}\right|+\tan ^{-1}\left(\frac{y}{x}\right)=c_1

\end{aligned}"

Which is the required isogonal trajectories of the given family of curves.